OpenStudy (anonymous):

At 10:00 a.m. two planes leave an airport. If the northbound plane flies at 280 mph and the southbound at 320 mph, at what time will they be 1,000 miles apart? Which equation can be used to solve for t, the time it takes the planes to be 1,000 miles apart?

OpenStudy (anonymous):

Let x be time. d=r*t Distance for the 1st plane is 280*x Distance for the 2nd plane is 320*x 280x+320x=1000 600x=1000 X=1000/600=5/3 h= 1h 40m Check: 280*5/3+320*5/3=1000 TRUE